OpenStudy (anonymous):

The system of equations is inconsistent. What are the missing values? 5x+__y=2 __x+3y=8

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

8 /2 = 4 5(4) = 20 3 / 4 = \(\frac{3}{4}\)

OpenStudy (anonymous):

so the top one is 4 and bottom is 20 ? @StudyGurl14

OpenStudy (studygurl14):

the top one is 3/4, the botom one is 20

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

it wont let me put a fraction in @StudyGurl14

OpenStudy (studygurl14):

It won't let you do 3/4 ?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

wait..i misread the question...hold on a sec

OpenStudy (studygurl14):

okay,so the top one is actually -3. to bottom is -5

OpenStudy (anonymous):

k thanks